OpenStudy (anonymous):

What are the solutions of 3x2 = 18x

Parth (parthkohli):

Divide both sides by \(3\) to start with.

Parth (parthkohli):

\( \color{Black}{\Rightarrow x^2 = 6x}\)

Parth (parthkohli):

Now take \(6x\) to the other side: \( \color{Black}{\Rightarrow x^2 - 6x = 0}\) \( \color{Black}{\Rightarrow x(x - 6) = 0}\)

Parth (parthkohli):

Solve for \(x\) in these equations: \( \color{Black}{\Rightarrow x = 0 \quad (\text{Already solved}) }\) \( \color{Black}{\Rightarrow x - 6 = 0}\)

OpenStudy (jiteshmeghwal9):

x(3x)=18x x=18x/3x=6x now take 6x to the other side x-6x=0 x(1-6)=0 x(-5)=0 x=5

Parth (parthkohli):

?

Parth (parthkohli):

\( \color{Black}{\Rightarrow 3 \times 5^2 = 18 \times 5}\) \( \color{Black}{\Rightarrow 75 = 90 }\) Equation does not satisfy for \(x = 5\).

OpenStudy (anonymous):

ohhh i see thanks